### v1

@@ 4.4.1 Pre-Saccade Potentials:
In this study, we prioritised the real-time nature of practical BCI tasks, which resulted in inter-saccade intervals too short to observe pre-saccade negativity or readiness potentials. The saccadic potentials of free-viewing saccades (see Figure 4(e-f)), revealed a signal that emerged approximately 50 ms before saccade onset, peaked around 50 ms after saccade execution, and lasted approximately from $[-50,200]$ ms relative to saccade onset. This timing aligns with the general description of saccade potentials, which typically begin within 100 ms before saccade onset and peak around the onset. The exact timing of pre-saccadic Spike Potentials (SPs) varies across studies. For example, Boylan et al. [56] reported that for horizontal saccades, SPs begin 12-4.4 ms before saccade onset, with peaks occurring between saccade onset and 7.8 ms post-saccade. Similarly, Doig et al. [57] found that for vertical saccades, SPs begin 17.3-6.8 ms before saccade onset, peaking between 3.4 ms pre-saccade and 5.2 ms post-saccade. Furthermore, Funase et al. demonstrated the existence of pre-saccade potentials across various saccade types, including visually-guided and auditory-guided saccades [58, 12], memory-guided saccades [59], and free-viewing saccades [15], with peak pre-saccade potentials occurring within 15 ms before saccade onset. Thus, the timing of saccade-related potentials observed for free-viewing saccades in our study is consistent with previous research.
@@ 4.4.2 Saccade Classification:
Although the results of the saccade onset classification reported in this study are high, the saccade direction classification does not outperform the results of the invasive oculomotor studies in non-human primates (discussed in the Introduction (Section 1)). The ECoG-based oculomotor BCI study [11] reported mean accuracies of 80% for two directions and 43% for four directions. For comparison, we evaluated the accuracy of our best-performing pipelines and achieved the following mean accuracies: saccade onset classification, 82% (chance 50%); four-class classification, 44% (chance 28%); left vs. up saccade, 67% (chance 52%); and left vs. down saccade, 76% (chance 55%). These values represent mean within-session decoding scores across three sessions (one session of Visually-Guided and two session of the Free-Viewing Task). The reported chance accuracies reflect the slight imbalances that existed in the number of trials for each direction, caused by the self-paced nature of the experiment and the removal of bad trials. Although the four-class decoding results are comparable, the ECoG study still outperformed our binary saccade direction classification results. However, considering that 80% is a minimum device accuracy benchmark for BCI usability [60], our system performed well for saccade onset classification. Furthermore, Ohmae et al. [9] reported similar findings, demonstrating that distinguishing saccades from fixations is a relatively easier task compared to classifying saccade directions. Their study showed that the best performance was achieved approximately 200 ms after saccade onset. This observation aligns with the saccadic potentials observed in IC7 for the Free-Viewing Task, which persist for about 200 ms after saccade onset, as illustrated in Figure 4(e-f).
@@ 4.4.3 Experiment Design Considerations:
The Visually-Guided and Free-Viewing Tasks differ in the neural features associated with saccade planning. Bruce et al. [52] demonstrated that a Visually-Guided Saccade Task activates three types of neurons: visual, visuomovement, and movement neurons. In contrast, free-viewing saccades occur without an external cue, engaging only movement neurons. This difference suggests that the Visually-Guided Task should elicit stronger neural signals, which may explain why previous invasive oculomotor studies predominantly used the Visually-Guided Task. For instance, Ohmae et al. [9] found that instead of relying solely on motor neurons, the performance of oculomotor BCI increased by incorporating all types of neurons that activate between stimulus presentation and saccade execution. Similarly, Tauscher et al. [14] compared the Visually-Guided and Free-Viewing Tasks in human participants, reporting more pronounced ERP responses time-locked to saccade onset in the Free-Viewing Task. However, confirming this distinction in our study is challenging, as the low-frequency ERP features of saccades overlap with the P300 response of visual cue in the Visually-Guided Task.
@@ 4.5 User-Centred and Translational Considerations for Home-Based BCI
Furthermore, this study was conducted at the participant’s home instead of a controlled laboratory setting, introducing practical and environmental challenges. During data collection, several uncontrolled factors were observed, such as noise from nearby construction and distractions from household pets. Although these factors increased variability and noise in the data, they also highlight the translational relevance and realism of the findings.
@@ 5 Conclusion
Appendix A
@@
Figure 10: Saccade classification performance using source space data. Each panel presents AUC-ROC scores for different decoding tasks: (a) initial saccades, (b) back saccades, and (c) combined initial and back saccades. Performance is reported for within-session (WS) decoding on the Visually-Guided Task and Free-Viewing Task (Sessions 1 and 2), as well as for cross-session (CS) decoding across sessions of the Free-Viewing Task. The red dashed line denotes chance-level performance (AUC = 0.5).

### v3

@@ 4.4.1 Pre-Saccade Potentials:
In this study, we prioritised the real-time nature of practical BCI tasks, which resulted in inter-saccade intervals too short to observe pre-saccade negativity or readiness potentials. The saccadic potentials of free-viewing saccades (see Figure 5(e-f)), revealed a signal that emerged approximately 50 ms before saccade onset, peaked around 50 ms after saccade execution, and lasted approximately from [50,200][-50,200] ms relative to saccade onset. This timing aligns with the general description of saccade potentials, which typically begin within 100 ms before saccade onset and peak around the onset. The exact timing of pre-saccadic Spike Potentials (SPs) varies across studies. For example, Boylan et al. [58] reported that, for horizontal saccades, SPs begin 12-4.4 ms before saccade onset, with peaks occurring between saccade onset and 7.8 ms post-saccade. Similarly, Doig et al. [59] found that, for vertical saccades, SPs begin 17.3-6.8 ms before saccade onset, peaking between 3.4 ms pre-saccade and 5.2 ms post-saccade. Furthermore, Funase et al. demonstrated the existence of pre-saccade potentials across various saccade types, including visually-guided and auditory-guided saccades [60, 13], memory-guided saccades [61], and free-viewing saccades [15], with peak pre-saccade potentials occurring within 15 ms before saccade onset. Thus, the timing of saccade-related potentials observed for free-viewing saccades in our study is consistent with previous research.
@@ 4.4.2 Saccade Classification:
Although the results of the saccade onset classification reported in this study are high, the saccade direction classification does not outperform the results of the invasive oculomotor studies in nonhuman primates (discussed in the Introduction (Section 1)). The ECoG-based oculomotor BCI study [8] reported mean accuracies of 80% for two directions and 43% for four directions. For comparison, we evaluated the accuracy of our best-performing pipelines and achieved the following mean accuracies: saccade onset classification, 82% (chance 50%); four-class classification, 44% (chance 28%); left vs. up saccade, 67% (chance 52%); and left vs. down saccade, 76% (chance 55%). These values represent mean within-session decoding scores across three sessions (one session of Visually-Guided and two sessions of the Free-Viewing Task). The reported chance accuracies reflect the slight imbalances that existed in the number of trials for each direction, caused by the self-paced nature of the experiment and the removal of bad trials. Although the four-class decoding results are comparable, the ECoG study still outperformed our binary saccade direction classification results. However, considering that 80% is a minimum device accuracy benchmark for BCI usability [62], our system performed well for saccade onset classification. Furthermore, Ohmae et al. [6] reported similar findings, demonstrating that distinguishing saccades from fixations is a relatively easier task compared to classifying saccade directions. Their study showed that the best performance was achieved approximately 200 ms after saccade onset. This observation aligns with the saccadic potentials observed in IC7 for the Free-Viewing Task, which persist for about 200 ms after saccade onset, as illustrated in Figure 5(e-f).
@@
Furthermore, cross-session decoding performance was lower compared to within-session classification. Decoding across sessions, particularly when data are collected on different days, is a more challenging BCI task due to neural signal non-stationarity, environmental variability, and changes in the user’s cognitive and physiological state. In our study, data collection was limited to two sessions for the Free-Viewing Task and one session for the Visually-Guided Task, due to the participant’s declining health. In contrast, developing a generalisable and robust cross-session classifier typically requires data from multiple sessions across different days. To mitigate this limitation, data augmentation techniques can be employed to artificially increase the effective size of the training set [63]. In addition to collecting more data in future studies, methods such as domain adaptation [64] and transfer learning [65] could also be used to improve generalisation across sessions.
@@ 4.4.3 Experiment Design Considerations:
The Visually-Guided Task and Free-Viewing Task differ in the neural features associated with saccade planning. Bruce et al. [66] demonstrated that a Visually-Guided Task activates three types of neurons: visual, visuomovement, and movement neurons. In contrast, free-viewing saccades occur without an external cue, engaging only movement neurons. This difference suggests that the Visually-Guided Task should elicit stronger neural signals, which may explain why previous invasive oculomotor studies predominantly used the Visually-Guided Task. For instance, Ohmae et al. [6] found that instead of relying solely on motor neurons, the performance of oculomotor BCI increased by incorporating all types of neurons that activate between stimulus presentation and saccade execution. Similarly, Tauscher et al. [14] compared the Visually-Guided and Free-Viewing Tasks in human participants, reporting more pronounced ERP responses time-locked to saccade onset in the Free-Viewing Task. However, confirming this distinction in our study is challenging, as the low-frequency ERP features of saccades overlap with the P300 response of visual cue in the Visually-Guided Task.
@@ 4.5 User-Centred and Translational Considerations for Home-Based BCI
Furthermore, this study was conducted at the participant’s home instead of a controlled laboratory setting, introducing practical and environmental challenges. During data collection, several uncontrolled factors were observed, such as noise from nearby construction and distractions from household pets. Although these factors increased variability and noise in the data, they also highlight the translational relevance of the findings.
@@ 5 Conclusion
Acknowledgment
This research was conducted in the Australian Research Council Training Centre in Cognitive Computing for Medical Technologies (Project Number ICI70200030), funded by the Australian Government. We thank the participant for providing consent to take part in this study and Z. Faraz for his assistance in collecting patient data. We thank all members of the ARC Training Centre and Synchron Inc. for their valuable feedback and discussions. We acknowledge the use of a large language model (ChatGPT, OpenAI) to assist in improving grammar and clarity, and the authors take full responsibility for the content of this manuscript.
Appendix A Comparison of Classification Algorithms
To select the most suitable classification algorithm, we compared the performance of five widely used methods: $k$-Nearest Neighbors (kNN), Logistic Regression, Linear Discriminant Analysis (LDA), Support Vector Machines (SVM) with an RBF kernel, and Random Forest. Figure 10 illustrates the classification performance of these algorithms using mean AUC-ROC scores. Each set of bar plots represents within-session (WS) and cross-session (CS) decoding performance for each classifier (x-axis). Panel (a) displays the results for saccade onset classification (fixation vs. saccade), while panels (b) and (c) show the performance for saccade direction classification tasks in four-class and binary (left vs. down) scenarios, respectively. The binary left vs. down comparison was specifically selected because it had demonstrated superior performance compared to other binary classification pairings in the main results.
Overall, performance across algorithms was comparable, with the Random Forest classifier performing better in certain cases, such as cross-session decoding for saccade onset classification (panel (a)). Since no single algorithm consistently outperformed others across all tasks, we calculated the average classification performance (mean $\pm$ standard deviation) for each algorithm. As shown in Table 2, Random Forest slightly outperformed other algorithms in both the saccade onset and binary direction (left vs. down) classification tasks.
Figure 10: Comparison of classification performance across different machine learning algorithms (x-axis) using sensor space data. Each panel shows AUC-ROC scores for a specific decoding task: (a) Saccade onset classification (fixation vs. saccade), (b) Four-class saccade direction classification, and (c) Binary direction classification (left vs. up saccade). Performance is reported for within-session (WS) decoding on both the Visually-Guided and Free-Viewing Tasks (Sessions 1 and 2), while cross-session (CS) decoding is reported for the Free-Viewing Task. The red dashed line indicates chance-level performance (AUC = 0.5).
Table 2: Classification performance (mean $\pm$ standard deviation of AUC-ROC scores) for five classifiers across three tasks. Scores are averaged across within-session and cross-session evaluations, and across both Visually-Guided and Free-Viewing paradigms. Bolded results indicate the best-performing classifier for each task.
Classification Algorithm Task
Saccade Onset Classification
Saccade Direction
Classification
4-Class
Binary
(Left vs. Down)
K Nearest Neighbours 0.81 ± 0.07 0.58 ± 0.05 0.67 ± 0.07
Logistic Regression 0.87 ± 0.05 0.63 ± 0.06 0.70 ± 0.08
Linear Discriminant Analysis 0.86 ± 0.07 0.63 ± 0.06 0.69 ± 0.08
Kernel SVM 0.86 ± 0.06 0.61 ± 0.03 0.66 ± 0.08
Random Forest 0.88 ± 0.03 0.63 ± 0.06 0.72 ± 0.07
Appendix B Saccadic Potentials
@@
Appendix C Decoding using Source Space Data
Figure 12: Saccade classification performance using source space data. Each panel presents AUC-ROC scores for different decoding tasks: (a) initial saccades, (b) back saccades, and (c) combined initial and back saccades. Performance is reported for within-session (WS) decoding on both the Visually-Guided and Free-Viewing Tasks (Sessions 1 and 2), while cross-session (CS) decoding is reported for the Free-Viewing Task. The red dashed line indicates chance-level performance (AUC = 0.5).